Operation Summary for February 25, 2011 - February 28, 2011 |
|
The Integrated Luminosity for the period from 2/21/11 to 2/28/11 was 41.81 pb-1 (inverse picobarns). NuMI reported receiving 3.98E18 protons on target during this same period. Friday Feb 25 The day shift began with Operations monitoring the Tevatron (TEV) store 8527 (28.94E30) and the Recycler stash (24.5E10), with the Antiproton Source (Pbar) stacking (15.1E10), with the Linac and Booster operating normally, with the Main Injector (MI) operating with MECAR problems, with NuMI off due to target LCW trouble, with the MiniBooNE experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations reported that MI was losing beam; Controls personnel changed a CAMAC card, which fixed the problem. TEV personnel conducted end-of-store studies from 10:35 AM to 2:44 PM. Operations terminated store 8527 at 2:57 PM and put the TeV into HEP shot setup at 3:09 PM. Cryo system technicians reported that they had tightened the belts on the Pbar dry engine. NuMI target LCW investigations continued. During the evening shift, Operations established store #8528 at 4:50 PM with an average initial luminosity (AIL) of 183.4E30. Operations reported that the Linac low energy Buncher went down at 10:25 PM. Linac personnel investigated the problem and began replacing the Buncher’s power amplifier tube (IPA2). Saturday Feb 26 The midnight shift began with Operations monitoring the store 8528 (72.08E30) and the stash (158.4E10), with Pbar stacking (22.8E10), and with MiniBooNE taking beam. During the midnight shift, Linac personnel reported that they had changed out the Buncher's IPA2 tube and socket by 1:39 AM and began tuning. At 3:45 AM, MI MECAR problems resulted in a loss of stacking. EE Support personnel investigated the MECAR problems for the rest of shift. During the day shift, EE Support personnel reported that the MECAR’s CPU bandwidth was overloaded and not functioning properly. They reinstalled the original MECAR hardware at MI-60, but this didn’t fix the problem; the fix was to load a 10-year-old memory image into the non-volatile memory and then load all MECAR parameters by hand. Pbar resumed stacking. Unfortunately, the MDAT frame delay offsets were set to incorrect values and caused BPM, MI beam permit, and V001 power supply issues. Pbar resumed stacking at 3:35 PM. During the evening shift, Operations terminated store 8528 at 5:11 PM, put the TEV into shot setup, and established store 8529 at 6:52 PM with an AIL of 247.58E30. NuMI personnel reported that they had removed some shielding blocks from the target enclosure and lowered a camera to the bottom; they saw no evidence of a LCW leak. They said the investigation would continue on Monday. Sunday Feb 27 The midnight shift began with Operations monitoring store 8529 and the stash, with Pbar stacking, and with MiniBooNE taking beam. During the midnight shift, Operations terminated store 8529 at 7:05 AM. Operations prepared D0 for an 8-hour access. During the day shift, Operations reported that D0 came out of access at 3:30 PM. Operators put the TEV through a low beta magnet checkout. Operators established a proton-only store and discovered a smoothing problem. Operators contacted TEV personnel for advice and then reloaded the dipole function generators ramp cards; the TEV sector C4 BPMs were in the wrong mode. During the evening shift, Operations put the TEV into shot setup at 7:39 PM and established store 8534 at 8:45 PM with an AIL of 337.7E30. Unfortunately, the store was lost due to a lightning strike at 9:52 PM. Monday Feb 28 The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE taking beam. During the midnight shift, Operations put the TEV into shot setup at 12:25 AM, but this potential store was also lost when a TEV RF station (TRF4) tripped off during the ramp and caused a three-house quench. Operations recovered the TEV and put it through a wet squeeze. The Plans for Monday The plans for today are to check the TEV RF and then resume normal stacking, stashing, and storing. |
|
More Information Comments and Suggestions |
| last modified 02/28/2011 email Fermilab |
| Security, Privacy, Legal | |